OpenStudy (anonymous):

Arrange these numbers from least to greatest: 2/5, 5/7, 4/9 please help

OpenStudy (anonymous):

ok, so first ur gonna want to turn the fractions into decimals

OpenStudy (anonymous):

do u know how 2 do that?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

ok, it really easy! just divide the numerator of the fraction by the denominator.

OpenStudy (anonymous):

can u do that and then post what u get? :)

OpenStudy (cj49):

2/5,=0.4 5/7=0.71 4/9=0.44

OpenStudy (texaschic101):

either that, or turn them into equivalent fractions with the same denominator...but I would suggest turning them all into decimals, it would be easier

OpenStudy (anonymous):

ok math confuses me all together

OpenStudy (anonymous):

ok, well @cj49 already posted the decimals.

OpenStudy (anonymous):

so i want u 2 put those decimals in order of least to greatest

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what did u get?

OpenStudy (anonymous):

0.4,0.44,0.71

OpenStudy (anonymous):

ok, so put the corresponding fractions in place of the decimals now :)

OpenStudy (midhun.madhu1987):